OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

\[\large\frac{ \frac{3}{4x+5}}{8}=4x\]??

OpenStudy (anonymous):

\[\frac{3}{4}x+\frac{5}{8}=4x\]?

OpenStudy (anonymous):

@satellite73 the second one you did